Dillon Labs has asked its financial manager to measure the cost of each specific type of capital as well as the weighted average cost of capital. The weighted average cost is to be measured by using the following​ weights: 50 % ​long-term debt, 25 % preferred​ stock, and 25 % common stock equity​ (retained earnings, new common​ stock, or​ both). The​ firm's tax rate is 23​%.

Debt: The firm can sell for ​$1010 a 16 ​-year, ​$1,000 ​-par-value bond paying annual interest at a 6.00 % coupon rate. A flotation cost of 2.5​% of the par value is required.

Preferred stock: 8.50 % ​(annual dividend) preferred stock having a par value of ​$100 can be sold for ​$94 . An additional fee of ​$3 per share must be paid to the underwriters.

Common stock: The​ firm's common stock is currently selling for ​$90 per share. The stock has paid a dividend that has gradually increased for many​ years, rising from ​$2.75 ten years ago to the ​$5.41 dividend​ payment, Upper D 0 ​, that the company just recently made. If the company wants to issue new new common​ stock, it will sell them ​$1.50 below the current market price to attract​ investors, and the company will pay ​$3.50 per share in flotation costs.  

a.  Calculate the​ after-tax cost of debt.

b.  Calculate the cost of preferred stock.

c.  Calculate the cost of common stock​ (both retained earnings and new common​ stock).

d.  Calculate the WACC for Dillon Labs.
Answer #1

a) After tax cost of debt (After tax YTM) :

YTM = (Coupon + ((F - P) / n)) / ((F + P) /2)

Here,

F (Face value) = $1,000

P (Net price) = Price - Flotation cost

P = $1,010 - ($1,000 * 2.5%)

P = $1,010 - $25 = $985

n (Years) = 16 years

Coupon = Face value * Coupon rate

Coupon = $1,000 * 6% = $60

Tax rate = 23% or 0.23

Now put the values into formula,

YTM = ($60 + (($1,000 - $985) / 16)) / (($1,000 + $985) / 2)

YTM = ($60 + $0.9375) / $992.50

YTM = 0.0614

After tax cost of debt = YTM * (1 - Tax rate)

After tax cost of debt = 0.0614 * (1 - 0.23)

After tax cost of debt = 0.0473 or 4.73%

b) Cost of preferred stock = Dividend / (Price - Flotation cost)

Here,

Dividend = Par value * Dividend rate

Dividend = $100 * 8.50% = $8.50

Price = $94

Flotation cost (Underwriters fee) = $3

Now,

Cost of preferred stock = $8.50 / ($94 - $3)

Cost of preferred stock = $8.50 / $91

Cost of preferred stock = 0.0934 or 9.34%

c) Cost of common stock :

i) Calculation of growth rate = ((End value / Start value) ^1/n) - 1

n (years) = 10 years

Growth rate = (($5.41 / $2.75)^1/10) - 1

Growth rate = ($1.97 ^ 1/10) - 1

Growth rate = 1.0702 (refer note) - 1

Growth rate = 0.0702 or 7.02%

Note : Rate calculation :

Step 1 : Take 1.97 on calculator & press √ sign 12 times

Step 2 : Deduct 1 & divide by 10 (ie. n)

Step 3 : Add back 1

Step 4 : Press multiply (*) and equals to (=) sign

Step 5 : Repeat step 4 for another 11 times

ii) Cost of common stock = D1 / (Price - Flotation cost) + g

Here,

g (growth rate) = 7.02% or 0.0702

D1 (Expected dividend) = Current Dividend + growth

D1 = $5.41 + ($5.41 * 7.02%)

D1 = $5.79

Price (net) = $90 - $1.50 (reduction) = $88.50

Flotation cost = $3.50

Now,

Cost of common stock = $5.79 / ($88.50 - $3.50) + 0.0702

Cost of common stock = ($5.79 / $85 ) + 0.0702

Cost of common stock = 0.0681 + 0.0702

Cost of common stock = 0.1383 or 13.83%

d) WACC = (Weight of debt * Cost of debt) + (Weight of preferred stock * Cost of preferred stock) + (Weight of common stock * Cost of common stock)

Here,

Weight of debt = 50% or 0.50

Weight of preferred stock = 25% or 0.25

Weight of common stock = 25% or 0.25

Cost of debt = 0.0473

Cost of preferred stock = 0.0934

Cost of common stock = 0.1383

Now,

WACC = (0.50 * 0.0473) + (0.25 * 0.0934) + (0.25 * 0.1383)

WACC = 0.0237 + 0.0234 + 0.0346

WACC = 0.0817 or 8.17%